OpenStudy (anonymous):

how do you verify cos(180 + x) = -cos(x)?

OpenStudy (watchmath):

You may use the formula \(cos(a+b)=\cos a\cos b-\sin a\sin b\). So \(cos(180+x)=\cos(180)\cos x -\sin(180)\sin x=(-1)\cos x - 0=-\cos x\)

OpenStudy (anonymous):

Thank you